There are 100g of CH4 and 100g of Cl2 gas in a 5 L flask at 35C. Calculate the partial pressure of each gas and the total pressure
1
Expert's answer
2020-05-15T10:22:46-0400

Assuming the ideal gas behavior of CH4 and Cl2, the partial pressure "p" can be calculated using the ideal gas law:

"p=\frac{nRT}{V}" ,

where "n" is the number of the moles of a gas, "R" is the ideal gas constant, 8.314 m3·Pa/(mol·K) , "T" is the temperature in kelvin, 35+273.15 = 308.15 K and "V" is the volume of the system, 5 L or 5·10-3 m3.

The number of the moles of a gas "n" is:

"n = \frac{m}{M}",

where "m" is the mass and "M" is the molar mass of a gas. The molar masses of CH4 and Cl2 are 16.04 g/mol and 70.91 g/mol, respectively. Therefore:

"n(CH_4) = \frac{100 \text{ g}}{16.04 \text{ g/mol}} = 6.23" mol,

"n(Cl_2) = \frac{100\text{ g}}{70.91\text{ g/mol}} = 1.41" mol.

The partial pressures of CH4 and Cl2 are:

"p_{CH_4} = \frac{6.23·8.314·308.15}{5·10^{-3}} = 3.19·10^6" Pa

"p_{Cl_2} = \frac{1.41·8.314·308.15}{5·10^{-3}} = 7.22·10^5" Pa.

According to the Dalton's law of partial pressures, the total pressure is the sum of the partial pressures:

"p_{tot} = p_{CH_4} + p_{Cl_2} = (3.19 + 0.722)\u00b710^6 = 3.91\u00b710^6" Pa.

Answer: the partial pressures of CH4 and Cl2 are 3.19·106 Pa and 7.22·105 Pa, respectively. The total pressure of the system is 3.91·106 Pa.